OpenStudy (callisto):

Given \(g(x,y) = \sqrt[3]{xy}\). Show that \(g_x(0,0)\) and \(g_y(0,0)\) exist by supplying values for them

OpenStudy (callisto):

If I am doing it right, then \[g_x= \frac{y}{3(xy)^\frac{2}{3}}\] If I put (x,y) = (0,0) into \(g_x\), then I’ll go into trouble

OpenStudy (anonymous):

This is a bizarre question.

OpenStudy (anonymous):

It's clearly not defined at those points. Do they mean to say the limit exists.

OpenStudy (callisto):

Part (a) of the question asks about the continuity of g(x,y), so I don't think it is asking if the limit exists.

OpenStudy (callisto):

And for the part after this, it actually asks if \(g_x\) and \(g_y\) are continuous at (0,0).

OpenStudy (anonymous):

Okay, here is the thing you have to understand: \(|x|\) is not differentiable at \(0\). The same thing applies \(\sqrt[2n]{x}\)

OpenStudy (anonymous):

Since \(\sqrt{}\) is the positive root.

OpenStudy (anonymous):

Okay how about this...

OpenStudy (anonymous):

\[ yy^{-2/3}=y^{1/3} \]Does that help?

OpenStudy (callisto):

But you still get 0/0

OpenStudy (anonymous):

You're not getting 0 out of that denominator as far as I can tell.

OpenStudy (callisto):

Why not? \[g_x= \frac{y}{3(xy)^\frac{2}{3}} = \frac{y^\frac{1}{3}}{3x^\frac{2}{3}}\]

OpenStudy (anonymous):

Try going back to the definition of a partial derivative maybe?

OpenStudy (anonymous):

I don't know how to go about this, but it seems wrong.

OpenStudy (callisto):

Do you mean I have to try finding \(g_x\) by the definition?

OpenStudy (anonymous):

The only thing I can think of is some loss of generalization that happened because or using differentiation rules.

OpenStudy (anonymous):

Do you need a refresh on the definition?

OpenStudy (anonymous):

\[ g_x(0,0) = \lim_{h\to 0}\frac{g(h,0)-g(0,0)}{h} \]

OpenStudy (anonymous):

That gives you \(0\)... woah.

OpenStudy (anonymous):

I guess using differentiation rules CAN screw you over.

OpenStudy (callisto):

Wait... \[g_x(0,0) = \lim_{h\to 0}\frac{g(h,0)-g(0,0)}{h} = \lim_{h\to 0}\frac{\sqrt[3]{h(0)}-\sqrt[3]{0\times 0}}{h} \]Right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

What is the issue?

OpenStudy (callisto):

\[g_x(0,0) = \lim_{h\to 0}\frac{g(h,0)-g(0,0)}{h} = \lim_{h\to 0}\frac{0}{h}\]Hmm..

OpenStudy (anonymous):

Did you have quote extension?

OpenStudy (anonymous):

Okay. \(0/h=0\)

OpenStudy (callisto):

Mind = blown Why applying differentiation rules doesn't work here? Or, where did I make a mistake?

OpenStudy (anonymous):

I have a couple of suspicions.

OpenStudy (callisto):

Yes?

OpenStudy (anonymous):

One is that the differentiation rules don't work the same way for partial derivatives.

OpenStudy (callisto):

But here's what Wolf gives us... http://www.wolframalpha.com/input/?i=partial+derivative+%5Csqrt%5B3%5D%7Bxy%7D

OpenStudy (anonymous):

The other is that there is an underlying assumption the differentiation rules are making which we are overlooking.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=partial+derivative+%5Csqrt%5B3%5D%7Bxy%7D+at+%280%2C0%29 The Wolf is not always perfect.

OpenStudy (anonymous):

I actually am trying to think of ways that this could happen with a single variable function, but I don't think it is possible.

OpenStudy (callisto):

At (0,0), it is not differentiable, what's the problem?

OpenStudy (anonymous):

It is differentiable by the definition though...

OpenStudy (callisto):

This one gives the general case, and if xy=0 http://www.wolframalpha.com/input/?i=partial+derivative+%5Csqrt%5B3%5D%7Bxy%7D This one gives a specific case, if x=y=0 http://www.wolframalpha.com/input/?i=partial+derivative+%5Csqrt%5B3%5D%7Bxy%7D+at+%280%2C0%29 I just don't understand how come is it differentiable, especially when you look at the graph..

OpenStudy (callisto):

* how come it is

OpenStudy (anonymous):

Because our concept of differentiable for single variable functions doesn't quite apply to multi variable functions the same way.

OpenStudy (anonymous):

You're only differentiating in a single direction.

OpenStudy (anonymous):

Okay, just consider: \[ g(x,0)=0 \]

OpenStudy (anonymous):

Clearly \(0\) is differentiable.

OpenStudy (callisto):

Are they really the same? Differentiate the function first, the evaluate it, and evaluate it first, then differentiate it.

OpenStudy (anonymous):

That is why it works. When you project it onto the \(xz\) plane it becomes differentiable. I think this phenomenon is exclusive to multi-variable derivatives. That makes it a uncommon phenomenon.

OpenStudy (anonymous):

@Callisto I'm not actually sure if they're the same by definition, but I think that the "projection" I'm talking about it part of the intent.

OpenStudy (anonymous):

the intent of partial derivatives.

OpenStudy (callisto):

Alright... So, to play safe, it is better to start with the definition, right?

OpenStudy (anonymous):

The definition is the true answer.

OpenStudy (anonymous):

Differentiation rules are a short cut.

OpenStudy (anonymous):

I think the insight to be taken from this is: If a multi-variable function returns an indeterminate form at a point, then going back to the original definition could provide some insight.

OpenStudy (anonymous):

That definition gives you the mathematical power of a limit. They are powerful devices.

OpenStudy (callisto):

Ok... Thanks for your help!!